OpenStudy (anonymous):

Rashida was paid $720 for 30 hours of work how much would it be for 40 hours of work? A 750 B 760 C 960 D 2880

OpenStudy (anonymous):

Id say A

OpenStudy (anonymous):

I agree with @abbychandler

OpenStudy (texaschic101):

proportion time 720/30 = x/40 cross multiply 30x = 28800 x = 28800/30 x = 960............not A...it is C

OpenStudy (texaschic101):

questions about anything I did ?